Oxidation number values of X , Y and Z of a compound are respectively +6 , +1 , and -2. Probable formula of this compound is Y2XZ4Explain plzz?????

Grade:11

1 Answers

SHAUNAK
7 years ago
Since the charge on the compound is zero, sum of charges of each element should be also zero.Therefore LHS:. X(1)+Y(2)+Z(4) =(+6)(1)+(+1)(2)+(-2)(4) =+6+2-8 =0.RHS=0.SINCE RHS=LHS, THE COMPOUND SHOULD HAVE MOLECULAR FORMULAX+2Y+4Z